OpenStudy (anonymous):

Answer this please fast, i'm confused. The local grocery store has videotapes on sale, $3.00 for 2 tapes. You have $20. a) How many tapes can you buy? b) If there is a 7% sales tax on the tapes, how many can you buy? (Explain your answers)

OpenStudy (anonymous):

a) 13 and b) 12

OpenStudy (anonymous):

How did you get the answer? explain please?

OpenStudy (anonymous):

\[\frac{20}{3} = 6\frac{2}{3}\] We times that by 2 to get the amount of tapes. This assumes you can buy the tapes in odd numbers and not forced to buy in pairs: \[2(\frac{20}{3})=\frac{40}{3}=13\frac{1}{3}\] We round down to 13. We times $3.00 by 1.07 to get the cost with sales. Then we repeat the process: \[3.00*1.07=3.21\] \[\frac{20}{3.21} \approx 6.23\] We again times by 2: \[2(6.23) = 12.46 \approx 12\]